# Machine Space I: Weak exponentials and quantification over compact spaces

### v3

@@ 3.2.1. The space of opens as a topological space
3.2.2. Pointfree quotients of machine space
For non-locally-compact locales, it is not obvious how to define the Scott topology on $\mathcal{O}X$ and the corresponding quotient machine space in a pointfree way. Nonetheless we can still prove some results that relate retracts of machine space to local compactness and existence of the exponential $\Sigma^{X}$. This section is more technical than previous ones and can be skipped by readers who are mainly interested in intuition.
We have seen that a section to the quotient $q\colon\Sigma^{\Sigma^{G}}\to\Sigma^{X}$ can be used to recover the evaluation map on $\Sigma^{X}$. In fact, we can say even more. Suplattice and preframe homomorphisms between frames are known to correspond to angelically and demonically nondeterministic maps respectively [24, 32, 31] and it would seem that a nondeterministic choice of representative for each open would be sufficient to construct the evaluation map. Even more generally we have the notion of triquotient maps.
Definition 3.14.
A locale map $f\colon X\to Y$ is a triquotient if there exists a dcpo morphism $f_{\#}\colon\mathcal{O}X\to\mathcal{O}Y$, called a triquotiency assignment, satisfying $f_{\#}(a\wedge f^{*}(b))=f_{\#}(a)\wedge b$ and $f_{\#}(a\vee f^{*}(b))=f_{\#}(a)\vee b$. (The map $f_{\#}$ is automatically a dcpo retraction of $f^{*}$.) ∎
Triquotients generalise both open and proper quotients and every retraction is a triquotient where we take the section as the triquotiency assignment. Plewe has shown in [19] that triquotients are stable under pullback and hence $q\times X$ is a triquotient whenever $q$ is. We now state a very general result about how to define the evaluation map (as long as $X$ is locally compact).
Proposition 3.15.
Let $X$ be a locally compact locale and suppose $q\colon\Sigma^{\Sigma^{G}}\to\Sigma^{X}$ is a triquotient. Then $\mathrm{ev}^{*}=t\circ\widetilde{\mathrm{ev}}^{*}$ where $t$ is any triquotiency assignment of $q\times X$.
Proof.
From the definition of $q$ we have the following diagram.
{tikzpicture}
[…]
Then $t\circ\widetilde{\mathrm{ev}}^{*}=t\circ(q^{*}\oplus\mathcal{O}X)\circ\mathrm{ev}^{*}=\mathrm{ev}^{*}$, as required. ∎
We now consider the situation when $X$ is not necessarily locally compact. Even without being able to easily define the space of opens in a pointfree way, we can still provide a kind of converse to Proposition 3.8.
For this we will need the definition of a continuous dcpo.
Definition 3.16.
The way-below relation $\ll$ on a dcpo $L$ is defined so that $a\ll b$ if and only if $b\leq\operatorname*{\bigvee{}^{\!\uparrow}}D$ implies there is some $d\in D$ with $a\leq d$. We say $L$ is continuous if for every $b\in L$ we have $b=\operatorname*{\bigvee{}^{\!\uparrow}}\{a\in L\mid a\ll b\}$. (Note, in particular, that the set $\{a\in L\mid a\ll b\}$ is required to be directed).
The result involves a general kind of map that generalises both retractions and open quotients and is incomparable with triquotients.
Definition 3.17.
We say a locale map $e\colon X\to Y$ is a semi-open quotient if $e^{*}$ has a suplattice retraction. ∎
Proposition 3.18.
If there is a semi-open quotient map $e\colon\Sigma^{\Sigma^{G}}\to\mathcal{O}X$ (where $\mathcal{O}X$ has the Scott topology) then $X$ is locally compact.
Proof.
Every locale of the form $\Sigma^{Y}$ is injective (see [13, Chapter VII]). Thus, in particular $\mathcal{O}(\Sigma^{\Sigma^{G}})$ is a projective frame and hence a retract of a free frame. But since $e$ is a semi-open quotient, $\mathcal{O}(\mathcal{O}X)$ is a suplattice retraction of $\mathcal{O}(\Sigma^{\Sigma^{G}})$. Thus, $\mathcal{O}(\mathcal{O}X)$ is a suplattice retraction of a free frame.
The free frame is constructed as a frame of downsets and is thus constructively completely distributive [9]. But a suplattice retract of a constructively completely distributive lattice is constructively completely distributive (see [20]) and hence so is $\mathcal{O}(\mathcal{O}X)$.
But it is shown in [27] that the points of a completely distributive frame are a continuous dcpo. Thus, $\mathcal{O}X$ is continuous and so $X$ is locally compact. ∎
If we could define the map $\overline{q}\colon\Sigma^{\Sigma^{G}}\to\mathcal{O}X$ and it were semi-open then this result would imply that $X$ is locally compact as in Proposition 3.12. It would be interesting to know if Proposition 3.18 also holds for triquotient maps.
@@ 3.2.2. Pointfree quotients of machine space
Another way to formulate these ideas is to embed $\mathrm{Loc}$ into the large presheaf category $\mathrm{Set}^{\mathrm{Loc}^{\mathrm{op}}}$, where exponentials of locales exist. Then if the map $\Sigma^{\Sigma^{G}}\to\Sigma^{X}$ induced by $X\hookrightarrow\Sigma^{G}$ has a section, it can be shown that $X$ is locally compact. In fact, the analogue of a dcpo section suffices. We omit the details.
@@ 4. Compactness and universal quantification
We believe that machine space also has the potential to be useful in clarifying other aspects of topology. In this section we use it to study compactness. Escardó has explained compactness in terms of an algorithm for universal quantification in the setting of semantics of programming languages [5, 7, 8]. Our approach yields a purely topological version of this algorithm.
@@ 4.1. An algorithm for universal quantification
It is important to understand exactly what it is such an algorithm needs to do. Of course, to give such a procedure we need to know a precise description of the space $X$ by generators and relations. From this we can mathematically derive whether a particular formal combination of generators is equal to $1$. On the other hand, recall that ‘the exact composition’ of the machines in machine space is opaque to us. Indeed, the opens of machine space only allow us to test very particular properties of the machines. In the first case we are dealing with discrete syntax, while in the second we are working with abstract machines which we do not have knowledge about a priori.
We are now ready to consider the algorithm for universal quantification over a compact locale $X$ with presentation $\langle G\mid R\rangle$. (To actually run such an algorithm we should restrict $G$ to be countable, but we can still imagine unbounded parallelism in theory.)
@@
     for each $S\in\mathcal{P}_{\mathsf{fin}}(\mathcal{P}_{\mathsf{fin}}(G))$ donormal-$\triangleright$ performed in parallel
@@
Note that the algorithm itself involves testing whether $\bigvee_{F\in S}\bigwedge_{g\in F}g$ is a cover of $X$. This is on the level of syntax and how to test this is determined mathematically before running the algorithm. The role of the algorithm is to translate from the level of syntax to the level of opaque machines.
@@
Let $X$ be a compact locale with $\mathcal{O}X=\langle G\mid R\rangle$. Algorithm 1 semi-decides if the given machine $m$ halts on all of $X$.
@@
The usual way to link compactness to universal quantification is via the characterisation in terms of closed product projections. Thus, it is perhaps remarkable that this algorithm instead makes use of the standard open cover definition.
@@
In 3.2 the procedure associated to ${\boxtimes}F$ is simply the act of observing the robot and comparing the generators it visits to the finite collection $F$. It is clear that this can be done uniformly in $F$. Alternatively, we might regard a machine as a program $m\colon\Sigma^{G}\to\Sigma$. To test whether a branch is contained in a subset $F$, we simply supply as input the generalised point $p_{F}\colon G\to\Sigma$ which halts precisely on the elements of $F$.
@@
Cantor space is the space $2^{\mathbb{N}}$ of infinite binary sequences. A presentation is given by $\langle z_{n},u_{n},\,n\in\mathbb{N}\mid z_{n}\wedge u_{n}=0,z_{n}\vee u_{n}=1\rangle$. Intuitively, $z_{n}$ should be thought of as the open consisting of the sequences whose $n^{\text{th}}$ digit is 00 and $u_{n}$ the open of sequences with $1$ in the $n^{\text{th}}$ position. A machine will be given by a formal expression of the form $\bigvee_{i\in I}\left(\bigwedge_{j\in J_{i}}z_{j}\wedge\bigwedge_{k\in K_{i}}u_{k}\right)$.
Note that a we can easily check if a finite join $\bigvee_{i\in I}\left(\bigwedge_{j\in J_{i}}z_{j}\wedge\bigwedge_{k\in K_{i}}u_{k}\right)$ is a cover of $2^{\mathbb{N}}$ by distributing the joins over the meets and checking each conjunct a cover. A conjunct will be a cover if and only if it contains both $z_{i}$ and $u_{i}$ as disjuncts for some $i\in\mathbb{N}$.
@@
We will explore how our approach relates to other representations of spaces and the link to Escardó’s setting in Section 4.3.
@@
[…]
Another element to discuss further is the process of determining whether $\bigvee_{F\in S}\bigwedge_{g\in F}g\sim 1$ with respect to $R$. We must do this by hand in order to construct the explicit algorithm for a given locale. In concrete cases this is typically straightforward (as it is for the examples above), but ideally we might hope to have a general algorithm that does this for us. However, generating a congruence from given relations can involve a transfinite procedure and is hence, at the very least, a computationally nontrivial process.
@@
Constructively, there is a ‘dual’ notion to compactness called overtness (see [26, 14]) which has the same relation to existential quantification as compactness has to universal quantification. One can give an algorithm for existential quantification that is very similar to Algorithm 1 except that instead of iterating over covers we iterate over ‘positive’ (classically: non-zero) elements. Classically, every locale is overt, though the resulting algorithm is still more subtle than naively searching through every real number in turn. Instead, it would use the fact that non-trivial open intervals with rational endpoints form a countable base for $\mathbb{R}$, which has the advantage of avoiding unbounded parallelism.
If the space $X$ is both compact and overt, we can combine the algorithms for universal and existential quantification to obtain an algorithm that determines whether a decidable predicate (i.e. clopen) holds everywhere on $X$ or not in finite time. This is most analogous to the case that is of primary interest to Escardó.
@@ 4.2. Topological consequences
It is possible to unwind Algorithm 1 to give an open of $\Sigma^{\Sigma^{G}}$. Each expression $\mathtt{test}(m\in{\boxtimes}F)$ is replaced with the corresponding open ${\boxtimes}F$. The second for loop takes the meet of these opens, $\bigwedge_{F\in S}{\boxtimes}F$. The first for (together with the if) corresponds to taking the join over all $S\in\mathcal{P}_{\mathsf{fin}}(\mathcal{P}_{\mathsf{fin}}(G))$ such that $\bigvee_{F\in S}\bigwedge_{g\in F}$ covers $X$. Let us denote the set of such $S$ by $\mathcal{C}$. Thus, the resulting open is $\bigvee_{S\in\mathcal{C}}\bigwedge_{F\in S}{\boxtimes}F$. When $X$ is compact, this is the open of machines which cover $X$. The following result is then immediate from Theorem 4.1.
@@
Let $X$ be a compact locale with $\mathcal{O}X=\langle G\mid R\rangle$ and let $\mathcal{C}$ the set $S\in\mathcal{P}_{\mathsf{fin}}(\mathcal{P}_{\mathsf{fin}}(G))$ satisfying $\bigvee_{F\in S}\bigwedge_{g\in F}g=1$ in $X$. Then the open $\bigvee_{S\in\mathcal{C}}\bigwedge_{F\in S}{\boxtimes}F$ contains precisely the machines which cover $X$.
@@
Also note Corollary 4.6 easily generalises to arbitrary compact sublocales $K$ of a general locale $X$ by simply viewing $K$ as direct sublocale of $\Sigma^{G}$.
@@
Corollary 4.6 also follows from the Hofmann–Mislove theorem [11]. Moreover, the Hofmann–Mislove theorem also provides a converse giving that the set of all covering machine of $X$ is Scott-open if and only if $X$ is compact. (See also [5, Lemma 7.4].) However, we believe restriction to machine space and the explicit algorithm provide an interesting new perspective on the situation.
@@
Of course, in the overt case we can similarly extract an open of all ‘positive’ machines for $X$ from the algorithm mentioned in Remark 4.5.
@@ 4.3. Link to domain-theoretic approaches
Recall that data types in programming languages can be modelled as certain kinds of dcpos. The dcpos that arise in denotational semantics are usually continuous, in which case the Scott topology is particularly well-behaved. We also note that resulting topological spaces are locally compact.
Escardó considers the data type of infinite binary sequences. We imagine these are implemented as streams where the terms are computed in order one by one on demand. Such a data consists not only of infinite sequences, but also additional elements corresponding to sequences whose first $n$ terms are defined, but for which accessing later terms in the sequence causes the computation to hang.
@@
In this way our construction is an example of a domain embedding where a topological space is represented as a subquotient of a dcpo. In the domain theory literature, there is some desire for spaces to occur as the subspace of maximal elements of a dcpo, but this not the case for our embedding, and indeed, this impossible if the space is not $T_{1}$. However, we can get closer to this ideal if we take the closure of $X$ as a subspace in $\mathcal{P}(G)$. A closed subset of a dcpo is a downward closed set which is closed under directed joins. Moreover, a closed subspace of a continuous dcpo is a continuous dcpo and the Scott topology on the subset agrees with the subspace topology. Thus, the closure of $X$ in $\mathcal{P}(G)$ is a dcpo which contains the points of $X$ and possibly some ‘more undefined’ approximations to them and since this closure is the smallest Scott-closed set containing $X$, the points of $X$ lie as close to the top of the dcpo as possible in some sense.
@@
Let us consider the case of Cantor space from Example 4.3. Recall that $\mathcal{O}(2^{\mathbb{N}})$ can be expressed as a quotient of $\mathcal{O}(\Sigma^{G})$ where $G=\{z_{n}\mid n\in\mathbb{N}\}\sqcup\{u_{n}\mid n\in\mathbb{N}\}$ by the congruence $C=\langle z_{n}\wedge u_{n}=0,z_{n}\vee u_{n}=1,\,n\in\mathbb{N}\rangle$. The closure of such a sublocale corresponds to the congruence generated by the pairs $(0,c)\in C$. In this case we obtain a the following presentation for the closure: $\langle z_{n},u_{n},\,n\in\mathbb{N}\mid z_{n}\wedge u_{n}=0\rangle$. This describes a dcpo the elements of which can be thought of as partial functions from $\mathbb{N}$ to $2=\{0,1\}$. The points of Cantor space correspond to the total functions, which do appear as the maximal elements of this dcpo. ∎
In general, the compactness algorithm works equally well if we replace $\mathcal{P}(G)$ with such a continuous dcpo $D$, or indeed any locally compact sublocale of machine space which contains $X$. The algorithm takes an element $m$ of $\Sigma^{D}$ as input and simple applies Algorithm 1 to $s(m)$ where $s$ is a section of the quotient $\Sigma^{\Sigma^{G}}\to\Sigma^{D}$ from Proposition 3.8.
@@
The inclusion $i\colon A\to\Sigma^{G}$ maps $f$ to a function which halts on $z_{n}$ if $f(n)=0$ and halts on $u_{n}$ if $f(n)=1$. The quotient map $q\colon\Sigma^{\Sigma^{G}}\to\Sigma^{A}$ sends $m$ to a map $x\mapsto m(i(x))$. Note that for a given $m\in\Sigma^{A}$, the above procedure $\mathtt{test}(s(m)\in{\boxtimes}F)$ for $F\in\mathcal{P}_{\mathsf{fin}}(G)$ can be easily extended to infinite $F\in\Sigma^{G}$. This gives the associated section $s\colon\Sigma^{A}\to\Sigma^{\Sigma^{G}}$ of $q$.
@@
$\mathcal{O}2^{\mathbb{N}}=\langle\ell_{p},\,p\in 2^{*}\mid{}$ $\ell_{p}\wedge\ell_{q}=\ell_{q}\text{ for $p\prec q$},$
$\ell_{p}\wedge\ell_{q}=0\text{ if $q\nprec p$ and $p\nprec q$},$
@@
where $p\prec q$ means $p$ is a prefix of $q$ and ##\## denotes concatenation.
However, taking the closure of this sublocale of $\Sigma^{2^{*}}$ does not yield the correct the dcpo. This is because it ignores the natural order structure on the prefixes. The solution is to modify the parent space $\Sigma^{G}$, so that instead of taking the free frame on a set $G$, we take the free frame on a poset. In this case, we equip $2^{*}$ with the reverse of the prefix order. Then $\mathcal{O}2^{\mathbb{N}}$ is the quotient of this frame by the congruence generated by the second and third relations above (since the first relation is already handled by the order structure on $G$). The closure of this sublocale then has the presentation $\langle\{\ell_{p}:p\in(2^{*})^{\mathrm{op}}\}\in\mathrm{Pos}\mid\ell_{p}\wedge\ell_{q}=0\text{ if $q\nprec p$ and $p\nprec q$}\rangle$. It is not hard to see that this is the locale of finite and infinite sequences binary sequences. (Indeed, it is the third relation in the presentation for $2^{\mathbb{N}}$ above that forces all sequences to be infinite.) This is precisely the dcpo considered by Escardó. (Finite sequences are interpreted as sequences which hang after some point.)
@@
Let us first consider how to check if a finite join $\bigvee_{i\in I}\left(\bigwedge_{p\in J_{i}}\ell_{p}\right)$ is a cover of $2^{\mathbb{N}}$. We first compute the finite meets using the first two relations to obtain a join of basic generators (omitting zeros from the join). Now let $N$ be the length of the longest prefix. We pad each $p$ in the join with 00s and $1$s to obtain all possible strings of length $N$ with prefix $p$. The join is a cover if and only if the resulting set contains every string of length $N$.
@@
The locale corresponding to free frame on a poset $P$ is $\Sigma^{\mathscr{I}P}$ where $\mathscr{I}P$ is the ideal completion of $P$. In the case above, the ideal completion of $(2^{*})^{\mathrm{op}}$ is actually already the desired dcpo. Nonetheless, it is still interesting to see that this fits into our general approach.

### v4

@@ 3.3. Relationship between machine space and the space of opens
Another way to formulate these ideas is to embed $\mathrm{Loc}$ into the large presheaf category $\mathrm{Set}^{\mathrm{Loc}^{\mathrm{op}}}$, where exponentials of locales exist. Then if the map $\Sigma^{\Sigma^{G}}\to\Sigma^{X}$ induced by $X\hookrightarrow\Sigma^{G}$ has a section, it can be shown that $X$ is locally compact. In fact, the analogue of a dcpo section (see [30]) suffices. We omit the details.
@@ 4. Compactness and universal quantification
We believe that machine space also has the potential to be useful in clarifying other aspects of topology. In this section we use it to study compactness. Escardó has explained compactness in terms of an algorithm for universal quantification in the setting of semantics of programming languages [7, 9, 10]. Our approach yields a version of this algorithm that involves only topological/localic concepts and the core operations provided by the verifiability interpretation, without restricting us to a particular programming language. Thus, we give a positive answer to second question posed in the introduction: there is a uniform algorithm for universal quantification over any compact space.
@@ 4.1. An algorithm for universal quantification
It is important to understand exactly what it is such an algorithm needs to do. Of course, to give such a procedure we need to know a precise description of the space $X$ by generators and relations. From this we can mathematically derive whether a particular formal combination of generators is equal to $1$. On the other hand, recall that ‘the exact composition’ of the machines in machine space is opaque to us. Indeed, the opens of machine space only allow us to test very particular properties of the machines. In the first case we are dealing with discrete ‘syntax’, while in the second we are working with abstract machines which we do not have knowledge about a priori (see Remark˜3.6).
We are now ready to consider the algorithm for universal quantification over a compact locale $X$ with presentation $\langle G\mid R\rangle$. (To actually run such an algorithm we should restrict $G$ to be countable, but we can still imagine uncountable parallelism in theory.)
@@
[…]
@@
Note that the algorithm itself involves testing whether $\bigvee_{F\in S}\bigwedge_{g\in F}g$ is a cover of $X$. This is on the ‘axiomatic’ level of Remark˜3.6 and how to test this is determined mathematically before running the algorithm. The role of the algorithm is to translate from the axiomatic level to the level of opaque machines.
@@
Let $X$ be a compact locale with $\mathcal{O}X=\langle G\mid R\rangle$. Algorithm˜1 semi-decides if the given machine $m$ halts on all of $X$.
@@
The usual way to link compactness to universal quantification is via the characterisation in terms of closed product projections (for example, see [3, 8, 18]). Thus, it is perhaps remarkable that this algorithm instead makes use of the standard open cover definition.
@@
In ˜3.7 the procedure associated to ${\boxtimes}F$ is simply the act of observing the machine and comparing the generators it visits to the finite collection $F$. It is clear that this can be done uniformly in $F$. Alternatively, we might regard a machine as a program $m\colon\Sigma^{G}\to\Sigma$. To test whether a branch is contained in a subset $F$, we simply supply as input the generalised point $p_{F}\colon G\to\Sigma$ which halts precisely on the elements of $F$.
@@
Cantor space is the space $2^{\mathbb{N}}$ of infinite binary sequences. A presentation is given by $\langle z_{n},u_{n},\,n\in\mathbb{N}\mid z_{n}\wedge u_{n}=0,z_{n}\vee u_{n}=1\rangle$. Intuitively, $z_{n}$ should be thought of as the open consisting of the sequences whose $n^{\text{th}}$ digit is 0 and $u_{n}$ the open of sequences with $1$ in the $n^{\text{th}}$ position. A machine will be given by a formal expression of the form $\bigvee_{i\in I}\left(\bigwedge_{j\in J_{i}}z_{j}\wedge\bigwedge_{k\in K_{i}}u_{k}\right)$.
Note that we can easily check if a finite join $\bigvee_{i\in I}\left(\bigwedge_{j\in J_{i}}z_{j}\wedge\bigwedge_{k\in K_{i}}u_{k}\right)$ is a cover of $2^{\mathbb{N}}$ by distributing the joins over the meets and checking that each conjunct is a cover. A conjunct will be a cover if and only if it contains both $z_{i}$ and $u_{i}$ as disjuncts for some $i\in\mathbb{N}$.
@@
We will explore how our approach relates to other representations of spaces and the link to Escardó’s setting in Section˜4.3.
@@
Consider the case of the closed real interval $X=[0,1]$. We can take a generating set consisting of certain intervals with rational end points, $G=\{[0,q)\mid q\in\mathbb{Q}\cap(0,1]\}\cup\{(q,1]\mid q\in\mathbb{Q}\cap[0,1)\}$. Then in the computational model, $[0,1]$ embeds as a subset of partial functions on $G$ and so a real $r$ corresponds to a function that halts on $[0,q)$ or $(q,1]$ if and only if $r\in[0,q)$ or $r\in(q,1]$ respectively. Then for a ‘machine’ $m\colon\Sigma^{G}\to\Sigma$ the algorithm halts if $m$ halts on all of the (functions corresponding to) reals in $[0,1]$. It is elementary to compute when a formal join of rational intervals covers $[0,1]$. Finally, $\mathtt{test}(m\in{\boxtimes}F\})$ is given by simply calling $m(p_{F})$ where $p_{F}\in\Sigma^{G}$ is given by $p_{F}(g)=\top\iff g\in F$. ∎
Another element to discuss further is the process of determining whether $\bigvee_{F\in S}\bigwedge_{g\in F}g\sim 1$ with respect to $R$. We must do this by hand in order to construct the explicit algorithm for a given locale. In concrete cases this is typically straightforward (as it is for the examples above), but ideally we might hope to have a general algorithm that does this for us. (In fact, it suffices for the algorithm to only semi-decide if the expression is a cover.) However, generating a congruence from given relations can involve a transfinite procedure and is hence, at the very least, a computationally nontrivial process.
@@
Constructively, there is a ‘dual’ notion to compactness called overtness (see [26, 15]) which has the same relation to existential quantification as compactness has to universal quantification. One can give an algorithm for existential quantification that is very similar to Algorithm˜1 except that instead of iterating over covers we iterate over ‘positive’ (classically: non-zero) elements. Classically, every locale is overt, though the resulting algorithm is still more subtle than naively searching through every real number in turn. Instead, it would use the fact that non-trivial open intervals with rational endpoints form a countable base for $\mathbb{R}$, which has the advantage of avoiding uncountable parallelism.
If the space $X$ is both compact and overt, we can combine the algorithms for universal and existential quantification to obtain an algorithm that determines whether a decidable predicate (i.e. clopen) holds everywhere on $X$ or not in finite time. This is analogous to the case that is of primary interest to Escardó.
@@ 4.2. Topological consequences
It is possible to unwind Algorithm˜1 to give an open of $\Sigma^{\Sigma^{G}}$. Each expression $\mathtt{test}(m\in{\boxtimes}F)$ is replaced with the corresponding open ${\boxtimes}F$. The second for loop takes the meet of these opens, $\bigwedge_{F\in S}{\boxtimes}F$. The first for (together with the if) corresponds to taking the join over all $S\in\mathcal{P}_{\mathsf{fin}}(\mathcal{P}_{\mathsf{fin}}(G))$ such that $\bigvee_{F\in S}\bigwedge_{g\in F}g$ covers $X$. Let us denote the set of such $S$ by $\mathcal{C}$. Thus, the resulting open is $\bigvee_{S\in\mathcal{C}}\bigwedge_{F\in S}{\boxtimes}F$. When $X$ is compact, this is the open of machines which cover $X$. The following result is then immediate from Theorem˜4.1.
@@
Let $X$ be a compact locale with $\mathcal{O}X=\langle G\mid R\rangle$ and let $\mathcal{C}$ be the set $S\in\mathcal{P}_{\mathsf{fin}}(\mathcal{P}_{\mathsf{fin}}(G))$ satisfying $\bigvee_{F\in S}\bigwedge_{g\in F}g=1$ in $X$. Then the open $\bigvee_{S\in\mathcal{C}}\bigwedge_{F\in S}{\boxtimes}F$ contains precisely the machines which cover $X$.
@@
Also note Corollary˜4.6 easily generalises to arbitrary compact sublocales $K$ of a general locale $X$ by simply viewing $K$ as direct sublocale of $\Sigma^{G}$.
@@
Corollary˜4.6 also follows from the Hofmann–Mislove theorem [12]. Moreover, the Hofmann–Mislove theorem also provides a converse giving that the set of all covering machine of $X$ is Scott-open if and only if $X$ is compact. (See also [7, Lemma 7.4].) However, we believe restriction to machine space and the explicit algorithm provide an interesting new perspective on the situation.
@@
Of course, in the overt case we can similarly extract an open of all ‘positive’ machines for $X$ from the algorithm mentioned in Remark˜4.5.
@@ 4.3. Link to domain-theoretic approaches
Recall that data types in programming languages can be modelled as certain kinds of dcpos. The dcpos that arise in denotational semantics are usually continuous, in which case the Scott topology is particularly well-behaved. We also note that the resulting topological spaces are locally compact.
Escardó considers the data type of infinite binary sequences. We imagine these are implemented as streams where the terms are computed in order one by one on demand. Such data consists not only of infinite sequences, but also additional elements corresponding to sequences whose first $n$ terms are defined, but for which accessing later terms in the sequence causes the computation to hang.
@@
In this way our construction is an example of a domain embedding where a topological space is represented as a subquotient of a dcpo. In the domain theory literature, there is some desire for spaces to occur as the subspace of maximal elements of a dcpo, but this not the case for our embedding, and indeed, this is impossible if the space is not $T_{1}$. However, we can get closer to this ideal if we take the closure of $X$ as a subspace in $\mathcal{P}(G)$. A closed subset of a dcpo is a downward closed set which is closed under directed joins. Moreover, a closed subspace of a continuous dcpo is a continuous dcpo and the Scott topology on the subset agrees with the subspace topology. Thus, the closure of $X$ in $\mathcal{P}(G)$ is a dcpo which contains the points of $X$ and possibly some ‘more undefined’ approximations to them and since this closure is the smallest Scott-closed set containing $X$, the points of $X$ lie as close to the top of the dcpo as possible in some sense.
@@
Let us consider the case of Cantor space from Example˜4.3. Recall that $\mathcal{O}(2^{\mathbb{N}})$ can be expressed as a quotient of $\mathcal{O}(\Sigma^{G})$ where $G=\{z_{n}\mid n\in\mathbb{N}\}\sqcup\{u_{n}\mid n\in\mathbb{N}\}$ by the congruence $C=\langle z_{n}\wedge u_{n}=0,z_{n}\vee u_{n}=1,\,n\in\mathbb{N}\rangle$. The closure of such a sublocale corresponds to the congruence generated by the pairs $(0,c)\in C$. In this case we obtain the following presentation for the closure: $\langle z_{n},u_{n},\,n\in\mathbb{N}\mid z_{n}\wedge u_{n}=0\rangle$. This describes a dcpo the elements of which can be thought of as partial functions from $\mathbb{N}$ to $2=\{0,1\}$. The points of Cantor space correspond to the total functions, which do appear as the maximal elements of this dcpo. ∎
In general, the compactness algorithm works equally well if we replace $\mathcal{P}(G)$ with such a continuous dcpo $D$, or indeed any locally compact sublocale of machine space which contains $X$. The algorithm takes an element $m$ of $\Sigma^{D}$ as input and simply applies Algorithm˜1 to $s(m)$ where $s$ is a section of the quotient $\Sigma^{\Sigma^{G}}\to\Sigma^{D}$ from Proposition˜3.10.
@@
The inclusion $i\colon A\to\Sigma^{G}$ maps $f\in A$ to a function which halts on $z_{n}$ if $f(n)=0$ and halts on $u_{n}$ if $f(n)=1$. The quotient map $q\colon\Sigma^{\Sigma^{G}}\to\Sigma^{A}$ sends $m$ to a map $x\mapsto m(i(x))$. Note that for a given $m\in\Sigma^{A}$, the above procedure $\mathtt{test}(s(m)\in{\boxtimes}F)$ for $F\in\mathcal{P}_{\mathsf{fin}}(G)$ can be easily extended to infinite $F\in\Sigma^{G}$. This gives the associated section $s\colon\Sigma^{A}\to\Sigma^{\Sigma^{G}}$ of $q$. ∎
@@
[…]
@@
where $p\prec q$ means $p$ is a prefix of $q$ and #\# denotes concatenation.
However, taking the closure of this sublocale of $\Sigma^{2^{*}}$ does not yield the correct dcpo. This is because it ignores the natural order structure on the prefixes. The solution is to modify the parent space $\Sigma^{G}$, so that instead of taking the free frame on a set $G$, we take the free frame on a poset. In this case, we equip $2^{*}$ with the reverse of the prefix order. Then $\mathcal{O}2^{\mathbb{N}}$ is the quotient of this frame by the congruence generated by the second and third relations above (since the first relation is already handled by the order structure on $G$). The closure of this sublocale then has the presentation $\langle\{\ell_{p}:p\in(2^{*})^{\mathrm{op}}\}\in\mathrm{Pos}\mid\ell_{p}\wedge\ell_{q}=0\text{ if $q\nprec p$ and $p\nprec q$}\rangle$. It is not hard to see that this is the locale of finite and infinite sequences binary sequences. (Indeed, it is the third relation in the presentation for $2^{\mathbb{N}}$ above that forces all sequences to be infinite.) This is precisely the dcpo considered by Escardó. (Finite sequences are interpreted as sequences which hang after some point.)
@@
Let us first consider how to check if a finite join $\bigvee_{i\in I}\left(\bigwedge_{p\in J_{i}}\ell_{p}\right)$ is a cover of $2^{\mathbb{N}}$. We first compute the finite meets using the first two relations to obtain a join of basic generators (omitting zeros from the join). Now let $N$ be the length of the longest prefix. We pad each $p$ in the join with 0s and $1$s to obtain all possible strings of length $N$ with prefix $p$. The join is a cover if and only if the resulting set contains every string of length $N$.
@@
The locale corresponding to the free frame on a poset $P$ is $\Sigma^{\mathscr{I}P}$ where $\mathscr{I}P$ is the ideal completion of $P$. In the case above, the ideal completion of $(2^{*})^{\mathrm{op}}$ is actually already the desired dcpo. Nonetheless, it is still interesting to see that this fits into our general approach.